# Learning to Drive Ethically: Embedding Moral Reasoning into Autonomous Driving

### v1

@@ 3.5 Cost function
$J_{Risk,ethic}=\omega_{B}J_{B}+\omega_{E}J_{E}+\omega_{M}J_{M},$ (11)
@@
$J_{Risk,ego}=\omega_{S}J_{S}$ (12)
with $J_{S}=\frac{\sum_{i=1}^{|S_{R}|}R_{i}}{|S_{R}|}$, but here, $R_{i}$ represents only the risk imposed on the ego vehicle by other participants. This formulation ignores the impact of the behavior of the ego vehicle on others and optimizes solely for self-preservation.
@@ 3.6 Prioritized experience replay
$\omega_{i}=(\frac{1}{NP_{i}})^{\beta},$
@@ 4.1 Simulation Environment
In this work, we use the MetaDrive simulator as the training and evaluation environment due to its ability to integrate real-world driving data, enabling the simulation of realistic driving scenarios (Li et al., 2022). This feature is crucial for our study, as we aim to assess the ethical decision-making capabilities of agents in settings that closely mirror real-world conditions, thereby enhancing the relevance and reliability of our evaluation. For the real-world data, we utilize the Waymo Open Dataset (Ettinger et al., 2021), which contains recordings from six cities across the United States. It includes a wide variety of challenging scenarios, such as unprotected turns, merges, lane changes, and complex intersection interactions, making it particularly well-suited for our task.
@@ 4.2 RL Setups
In our designed scenarios, we focus on ethical decision-making and assume the existence of a reliable perception module. This module provides accurate, real-time information about surrounding traffic participants, including vehicles, pedestrians, and cyclists. As a result, our RL agent is dedicated solely to decision-making rather than perception. To support this task, we define the observation of the agent at each timestep $t$ as $o_{t}$. We use the Frenet coordinate system instead of Cartesian coordinates for both the observation and action spaces. The Frenet system describes motion relative to a reference path using two parameters: s, the longitudinal distance along the path, and d, the lateral offset from the path. This representation simplifies trajectory planning on curved roads by decoupling progress and deviation, providing an intuitive view of the vehicle’s relationship to the road. The observation $o_{t}$ is a stacked vector composed of four components: $o_{ov,t}$, containing features of the ego vehicle; $o_{on,t}$, representing navigation-related features; $o_{sv,t}$, describing the surrounding vehicles; and $o_{sp,t}$, which accounts for other traffic participants such as pedestrians and cyclists.
@@
$o_{ov,t}=\left(\frac{d_{l,t}}{d_{w}},\frac{d_{r,t}}{d_{w}},\frac{v_{ev,t}}{v_{max}},\frac{[\psi_{ev,t}-\chi_{ev,t}]_{-\pi}^{\pi}}{\pi},\dot{\psi}_{ev,t}\right),$ (14)
where $d_{l,t}$ and $d_{r,t}$ are the lateral distances from the ego vehicle to the left and right road boundaries, normalized by the road width $d_{w}$. The term $v_{ev,t}$ is the ego speed, normalized by the maximum allowable speed $v_{max}$. The heading angle $\psi_{ev,t}$ and yaw rate $\dot{\psi}{ev,t}$ describe vehicle orientation, while $\chi{ev,t}$ is the heading of the road at the location of the ego vehicle.
@@
$o_{on,t}^{wp}=\left(\frac{\Delta d_{d,t}}{d_{scale}},\frac{\Delta d_{s,t}}{d_{scale}},\frac{\kappa_{l,t}}{d_{norm}},\frac{\psi_{l,t}}{\psi_{norm}}\right),$ (15)
@@
Since the objective of the agent is path planning, it must generate a smooth polynomial trajectory in the Frenet space at each timestep $t$. The action $a_{t}$ consists of three continuous values: $a_{t}=\left(T_{t},D_{d,t},v_{s,t}\right)$, with $T_{t}$ indicates the target duration of the trajectory, $D_{d,t}$ and $v_{s,t}$ describe the target lateral position and target longitudinal speed along the reference path. To ensure the stability of the generated trajectories, we constrained the action with $T_{t}\in[0,2]$, $D_{d,t}\in[0,d_{w}/2]$, and $v_{s,t}\in[0,v_{max}]$. The actions are normalized during training to enhance learning stability. Given these parameters, the agent generates a polynomial trajectory using the path planner. More details are provided in Section 4.3.
@@
[…]
$\Delta R_{\text{terminal}}=\left\{\begin{array}[]{ l l r }R_{\text{success}}&\text{Arrival at destination}\\ R_{\text{out}}&\text{ Out of road}\\ R_{\text{collision}}&\text{ Collision}\end{array}\right$. (19)
Table 1 summarizes the reward components and their corresponding functions. The hyperparameters used in this reward formulation are listed in Table 5. Overall, this reward function encourages safe, efficient, and goal-oriented behavior. It balances competing objectives such as maintaining speed, staying on track, and avoiding abrupt maneuvers or collisions. The modular structure also allows easy adjustment of weights and terms for different driving scenarios and training goals.
@@ 4.3.1 Path Planner
To ensure the comfort of the generated trajectory $\xi_{\tau,t}$ at time step $t$, a quartic polynomial in the longitudinal direction $s$ and a quintic polynomial in the lateral direction $d$ are employed, such that, $\xi_{\tau,t}=\{s(\tau),d(\tau)\}$. This formulation follows established practice, as it guarantees the continuity of both longitudinal and lateral accelerations, thereby contributing to a smooth and comfortable trajectory.
$\left\{\begin{array}[]{ l l }s(\tau)&{}=a_{0}+a_{1}\tau+a_{2}\tau^{2}+a_{3}\tau^{3},+a_{4}\tau^{4}\\ d(\tau)&{}=b_{0}+b_{1}\tau+b_{2}\tau^{2}+b_{3}\tau^{3},+b_{4}\tau^{4}+b_{5}\tau^{5},\end{array}\right$. (20)
for $\tau\in\{0,\Delta\tau,2\Delta\tau,...,T\}$, where $\{a_{1},...,a_{4}\}$ and $\{b_{1},...,b_{5}\}$ denote the polynomial coefficients in the longitudinal and lateral directions, respectively. To solve equation (20), both the initial vehicle state $s_{0},\dot{s}_{0},\ddot{s}_{0},d_{0},\dot{d}_{0},\ddot{d}_{0}]$ and the terminal state at the planning horizon $T$, given by, $\dot{s}_{T},\ddot{s}_{T},d_{T},\dot{d}_{T},\ddot{d}_{T}]$, must be specified.
Following the approach in (Werling et al., 2010), and under the assumptions $\ddot{s}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort. In this framework, vehicle trajectories are generated to achieve a target longitudinal speed $\dot{s}_{T}$ and a desired lateral position $d_{T}$. To enable the RL agent to adaptively control driving behavior in response to dynamic traffic environments, the planning horizon $T$ is treated as an action output of the agent. This design allows the agent to flexibly determine the duration of each trajectory segment. Subsequently, a sequence of waypoints is produced using equation (20), which are then tracked by a path-following controller. To discourage the generation of uncomfortable or impractical trajectories, a penalty term on longitudinal jerk is incorporated into the reward function, as detailed in Section 4.2. This encourages the agent to generate smooth and context-appropriate trajectories.
@@ 5.2 Worst-case analyses
After evaluating the general performance of the different agents, we proceed to assess their behavior in worst-case scenarios. For this analysis, we use the Time-to-Collision (TTC) metric, which estimates the time remaining before a collision would occur between traffic participants if they continue at their current speeds and along their current paths. TTC is a critical indicator for assessing safety. To capture these scenarios, we log the minimum TTC between the ego vehicle and other vehicles, along with the corresponding maximum other risk values at each timestep during evaluation. The rationale is that a predictive and cautious agent should be able to reduce risk in advance, resulting in fewer instances where both TTC is low and other risk is high. Conversely, a higher frequency of such cases indicates a failure to anticipate and mitigate danger. The results are visualized in Figure 5, using a cost limit of 0.6 for all agents. For clarity, we define undesirable behavior as instances where TTC is less than 2 seconds and other risk exceeds 0.25. These critical cases are highlighted with red squares in the figure.
@@
In Figure 5 (a), we compare the EthicAR agent in both ethical and selfish modes, using the standard mode as a baseline. As shown in the heatmap, the standard mode exhibits the highest number of undesirable behaviors, specifically, instances with high risk and low TTC, indicating that it fails to mitigate risk and is trained solely based on the reward function defined in Equation 17. In contrast, both the ethical and selfish modes of EthicAR demonstrate significantly fewer high-risk, low-TTC cases, with the ethical mode performing best. This suggests that incorporating ethical considerations, even in the selfish mode, leads to safer behavior compared to the standard mode. Figure 5 (b) presents a cross-agent comparison, showing the ethical mode results for three agents: SACLAG, EthicAR, and EthicAR w/o PER. All three agents are able to reduce the frequency of undesirable behaviors compared to the standard mode. Among them, EthicAR achieves the best performance, further highlighting the effectiveness of its design.
@@ 5.3 Scenario analyses
The standard-mode agent, trained solely on efficiency and safety objectives with the reward function (17), exhibits its limitations in this scenario. Starting from a TTC to the cyclist of 3.2s and a low-risk situation, the standard mode, like other baselines, adjusts its heading to achieve a TTC of inf by laterally diverging from the cyclist. However, with no ethical constraints, it immediately attempts to overtake once the TTC reaches inf, closing to an uncomfortably small lateral gap. In reality, forcing an overtake on a narrow, two-lane road with oncoming congestion violates both ethical norms and common etiquette (Luetge, 2017), and risks a collision either with the cyclist or opposing traffic due to the unpredictability of the cyclist. Such unsafe aggressiveness cannot be mitigated by other baselines: both the PER-ablated EthicAR and the SACLAG agent display similarly risky overtaking attempts here. By contrast, EthicAR maintains a respectful distance from the cyclist throughout and refrains from overtaking, keeping its maximum other risk at a low 0.1 until the conclusion of the scenario, demonstrating more socially acceptable, ethically aligned behavior suitable for everyday driving.
@@
In everyday driving etiquette at an unsignalized intersection, especially when the ego vehicle approaches from the right, it technically has priority to proceed first. However, safety considerations lead all agents to active yield. While this reflects legal norms, not all agents minimize risk effectively. For instance, for the standard mode, starting from inf TTC, and risk of 0, and then allows the other vehicle to turn. Yet its approach remains aggressive: by T=4s, TTC falls to just 0.4s and risk spikes to 0.7, even though no collision occurs. Such behavior is risky and unsuitable for AV deployment. EthicAR w/o PER agent displays a similar pattern. The SACLAG agent demonstrates improved restraint: though ego speed remains relatively high at T=4s, TTC and risk increase to about 0.3, still yielding more courteously. By contrast, EthicAR maintains the lowest risk level throughout the scenario, embodying responsible and socially considerate driving.
@@ 6 Discussion
In this work, we introduce EthicAR, a hierarchical autonomous driving framework designed to minimize the overall risk to all road users and make ethically responsible decisions. EthicAR does not only react ethically in high-risk situations but also incorporates ethical considerations into routine driving behavior, thereby proactively avoiding dangerous scenarios. The framework consists of two levels: a decision level and an execution level. The decision level is powered by an LSTM-based SACLag algorithm, enhanced with dynamic PER and trained under ethical cost constraints. The execution level ensures that planned actions are translated into smooth and comfortable vehicle control. To evaluate the contributions of each component, we conduct an ablation study examining the effects of PER, the LSTM architecture, and different cost function designs. The agents are trained using a real-world dataset in simulation, providing realistic scenarios that improve generalization and reliability. For performance evaluation, we test the trained agents on 75 previously unseen real-world scenarios. We analyze the resulting risks for both the ego vehicle and surrounding road users. Our results show that the ethical mode of EthicAR outperforms the other variants, significantly reducing overall risk for all road users compared to versions without cost constraints. Interestingly, although EthicAR is explicitly optimized to reduce risk to all road users, it also achieves lower ego risk than the selfish mode, which considers only the safety of the ego vehicle. This suggests that ethical behavior, inherently designed to protect others, also benefits the ego vehicle, due to the interactive nature of traffic dynamics. Additionally, we analyze four real-world scenarios involving ethical dilemmas. These case studies confirm that EthicAR consistently makes responsible decisions, not only toward other vehicles but also toward vulnerable road users such as cyclists. Looking ahead, extending the framework to incorporate trajectory predictions for pedestrians and cyclists, in addition to vehicles, will enable even more realistic behavior modeling and improve readiness for real-world deployment.

### v2

@@ 3.5 Cost function
$J_{\text{ethic}}=\omega_{B}J_{B}+\omega_{E}J_{E}+\omega_{M}J_{M},$ (11)
@@
$J_{\text{ego}}=\omega_{S}J_{S}$ (12)
with $J_{S}=\frac{\sum_{i=1}^{|S_{R}|}R_{i,e}}{|S_{R}|}$. Here, $R_{i,e}$ denotes the risk imposed on the ego vehicle by other participants. Although the form of $J_{S}$ resembles that of $J_{B}$, the two quantify different objectives. While $J_{B}$ accounts for the risks experienced by all traffic participants, $J_{S}$ ignores the effects of the ego vehicle on others and focuses solely on minimizing its own risk, effectively optimizing for self-preservation.
To further protect VRUs, the current cost function can be extended to incorporate preferences for lower VRU risk within its three components. In this paper, we adopt an equal treatment of all road users. However, prioritization of VRUs can be achieved by introducing weighting factors to the risk and harm terms associated with them, thereby placing greater emphasis on their protection.
@@ 3.6 Prioritized experience replay
$\omega_{i}=\frac{1}{{(NP_{i})}^{\beta}},$
@@ 4.1 Simulation Environment
In this work, we use the MetaDrive simulator (Li et al., 2022) as the training and evaluation environment due to its ability to integrate real-world driving data, enabling the simulation of realistic driving scenarios. This feature is crucial for our study, as we aim to assess the ethical decision-making capabilities of agents in settings that closely mirror real-world conditions, thereby enhancing the relevance and reliability of our evaluation. For the real-world data, we utilize the Waymo Open Dataset (Ettinger et al., 2021), which contains recordings from six cities across the United States. It includes a wide variety of challenging scenarios, such as unprotected turns, merges, lane changes, and complex intersection interactions, making it particularly well-suited for our task.
@@ 4.2 RL Setups
In our designed scenarios, we focus on ethical decision-making and assume the existence of a reliable perception module. This module provides accurate, real-time information about surrounding traffic participants, including vehicles, pedestrians, and cyclists. As a result, our RL agent is dedicated solely to decision-making rather than perception. To support this task, we define the observation of the agent at each timestep $t$ as $o_{t}$. We use the Frenet coordinate system instead of Cartesian coordinates for both the observation and action spaces. The observation $o_{t}$ is a stacked vector composed of four components: $o_{ov,t}$, containing features of the ego vehicle; $o_{on,t}$, representing navigation-related features; $o_{sv,t}$, describing the surrounding vehicles; and $o_{sp,t}$, which accounts for other traffic participants such as pedestrians and cyclists.
@@
$o_{ov,t}=\left(\frac{d_{l,t}}{d_{w}},\frac{d_{r,t}}{d_{w}},\frac{v_{ev,t}}{v_{max}},\frac{[\psi_{ev,t}-\chi_{ev,t}]_{-\pi}^{\pi}}{\pi},\dot{\psi}_{ev,t}\right)^{\top},$ (14)
where $d_{l,t}$ and $d_{r,t}$ are the lateral distances from the ego vehicle to the left and right road boundaries, normalized by the road width $d_{w}$. The term $v_{ev,t}$ is the ego speed, normalized by the maximum allowable speed $v_{max}$. The heading angle $\psi_{ev,t}$ and yaw rate $\dot{\psi}_{ev,t}$ describe vehicle orientation, while $\chi_{ev,t}$ is the heading of the road at the location of the ego vehicle.
@@
$o_{on,t}^{wp}=\left(\frac{\Delta d_{d,t}}{d_{scale}},\frac{\Delta d_{s,t}}{d_{scale}},\frac{\kappa_{l,t}}{d_{norm}},\frac{\psi_{l,t}}{\psi_{norm}}\right)^{\top},$ (15)
@@
Since the objective of the agent is path planning, it must generate a smooth polynomial trajectory in the Frenet space at each timestep $t$. As discussed in Section 3.1, the action $a_{t}$ is defined as $a_{t}=\left(T_{t},D_{d,t},v_{s,t}\right)$. To ensure the stability of the generated trajectories, we constrained the action with $T_{t}\in[0,2]$, $D_{d,t}\in[0,d_{w}/2]$, and $v_{s,t}\in[0,v_{max}]$. The actions are normalized during training to enhance learning stability. Given these parameters, the agent generates a polynomial trajectory using the path planner. More details are provided in Section 4.3.
@@
with three components, the speed reward $r_{v,\tau}$, the progress reward $r_{p,\tau}$, and a one-time terminal reward $\Delta R_{\text{terminal}}$. The weights $\omega_{v}$ and $\omega_{p}$ control the importance of the speed and progress terms. The speed reward encourages the agent to maintain a desired speed $v_{\text{des}}$. It uses a binary direction flag $\rho_{pr}\in\{-1,+1\}$, which indicates whether the vehicle is moving in the correct direction. This helps enforce directionality and penalizes driving the wrong way. The progress reward gives the agent a small positive reward $\Delta s$ for moving forward in the correct direction. This incentivizes consistent progress along the path. The terminal reward $\Delta R_{\text{terminal}}$ is given only when the agent reaches a terminal condition:
$\Delta R_{\text{terminal}}=\left\{\begin{array}[]{ll}R_{\text{success}}&\quad\text{Arrival at destination},\\[2.0pt] R_{\text{out}}&\quad\text{Out of road},\\[2.0pt] R_{\text{collision}}&\quad\text{Collision},\end{array}\right$. (19)
where the success reward is set to $R_{\text{success}}=25$, the off-road penalty to $R_{\text{out}}=-15$, and the collision penalty depends on the type of object involved: $R_{\text{collision}}=-10$ for collisions with other vehicles and $R_{\text{collision}}=-15$ for collisions with vulnerable road users such as pedestrians or cyclists. Table 1 summarizes the reward components and their corresponding functions. The hyperparameters used in this reward formulation are listed in Table 5. Overall, this reward function encourages safe, efficient, and goal-oriented behavior. It balances competing objectives such as maintaining speed, staying on track, and avoiding abrupt maneuvers or collisions. The modular structure also allows easy adjustment of weights and terms for different driving scenarios and training goals.
@@ 4.3.1 Path Planner
To ensure the comfort of the generated trajectory $\xi_{e,t}$ at time step $t$, a quartic polynomial in the longitudinal direction $l$ and a quintic polynomial in the lateral direction $d$ are employed, such that, $\xi_{e,t}=\{l_{e,\tau},d_{e,\tau}\}$. This formulation follows established practice, as it guarantees the continuity of both longitudinal and lateral accelerations, thereby contributing to a smooth and comfortable trajectory.
$\left\{\begin{array}[]{ l l }l_{e,\tau}&{}=a_{0}+a_{1}\tau+a_{2}\tau^{2}+a_{3}\tau^{3}+a_{4}\tau^{4},\\ d_{e,\tau}&{}=b_{0}+b_{1}\tau+b_{2}\tau^{2}+b_{3}\tau^{3}+b_{4}\tau^{4}+b_{5}\tau^{5},\end{array}\right$. (20)
for $\tau\in\{0,\Delta\tau,2\Delta\tau,...,T\}$, where $\{a_{1},...,a_{4}\}$ and $\{b_{1},...,b_{5}\}$ denote the polynomial coefficients in the longitudinal and lateral directions, respectively. To solve equation (20), both the initial vehicle state $l_{0},\dot{l}_{0},\ddot{l}_{0},d_{0},\dot{d}_{0},\ddot{d}_{0}]$ and the terminal state at the planning horizon $T$, given by, $\dot{l}_{T},\ddot{l}_{T},d_{T},\dot{d}_{T},\ddot{d}_{T}]$, must be specified.
Following the approach in (Werling et al., 2010), and under the assumptions $\ddot{l}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort. In this framework, vehicle trajectories are generated to achieve a target longitudinal speed $\dot{l}_{T}$ and a desired lateral position $d_{T}$. To enable the RL agent to adaptively control driving behavior in response to dynamic traffic environments, the planning horizon $T$ is treated as an action output of the agent. This design allows the agent to flexibly determine the duration of each trajectory segment. With the solutions appended in Section A.2 for equation (20), a sequence of waypoints is produced, which are then tracked by a path-following controller. To discourage the generation of uncomfortable or impractical trajectories, a penalty term on longitudinal jerk is incorporated into the reward function, as detailed in Section 4.2. This encourages the agent to generate smooth and context-appropriate trajectories.
@@ 5.2 Worst-case analyses
After evaluating the general performance of the different agents, we proceed to assess their behavior in worst-case scenarios. For this purpose, we use the Time-to-Collision (TTC) metric, which estimates the time remaining before a collision would occur between traffic participants if they continue at their current speeds and along their current paths. TTC is a critical indicator for assessing safety. To capture these scenarios, we log the minimum TTC between the ego vehicle and other vehicles, together with the corresponding maximum other risk values at each timestep during evaluation. The underlying rationale is that a predictive and cautious agent should be able to reduce risk in advance, leading to fewer instances where both TTC is low and other risk is high. Conversely, a higher frequency of such cases indicates a failure to anticipate and mitigate danger. The results are visualized in Figure 5, using a cost limit of 0.6 for all agents. The heatmap encodes the frequency of TTC–risk pairs: darker cells correspond to higher frequencies. For clarity, we define undesirable behavior as cases where TTC is less than 2 seconds and other risk exceeds 0.25. These critical cases are highlighted with red squares.
@@
In Figure 5 (a), we compare the EthicAR agent in both ethical and selfish modes against the standard mode baseline. Within the critical area, the heatmap shows that the standard mode has the most darker cells, indicating the largest number of high-risk, low-TTC cases. This highlights its inability to mitigate danger when trained solely on the reward function (17). In contrast, both the ethical and selfish modes of EthicAR show markedly fewer critical cases, with the ethical mode performing best: only three isolated instances fall into the red squares. This suggests that incorporating ethical considerations, even in the selfish mode, leads to safer behavior compared to the standard mode. Figure 5(b) extends the comparison across agents, focusing on the ethical mode of SACLAG, EthicAR, and EthicAR w/o PER. All three agents reduce the frequency of undesirable behaviors relative to the standard mode. Among them, EthicAR achieves the strongest performance, further highlighting the effectiveness of its design.
@@ 5.3 Scenario analyses
The standard-mode agent, trained solely on efficiency and safety objectives with the reward function (17), exhibits its limitations in this scenario. Starting from a TTC to the cyclist of 3.2s and a low-risk situation, the standard mode, like other baselines, adjusts its heading to achieve a TTC of inf by laterally diverging from the cyclist. However, with no ethical constraints, it immediately attempts to overtake once the TTC reaches inf, closing to an uncomfortably small lateral gap. In reality, forcing an overtake on a narrow, two-lane road with oncoming congestion violates both ethical norms and common etiquette (Luetge, 2017), and risks a collision either with the cyclist or opposing traffic due to the unpredictability of the cyclist. Such unsafe aggressiveness cannot be mitigated by other baselines: both the PER-ablated EthicAR and the SACLAG agent display similarly risky overtaking attempts here. By contrast, EthicAR maintains a respectful distance from the cyclist throughout and refrains from overtaking, keeping its maximum other risk at a low 0.1 until the conclusion of the scenario, demonstrating more socially acceptable, ethically aligned behavior suitable for everyday driving. Importantly, despite the cost function not explicitly prioritizing the minimization of risks to VRUs, EthicAR nonetheless exhibited responsible decision-making toward them. This outcome underscores the effectiveness of the proposed framework and further suggests that the protection of VRUs could be enhanced through straightforward adjustments to weight assignments.
@@
In everyday driving etiquette at an unsignalized intersection, especially when the ego vehicle approaches from the right, it technically has priority to proceed first. However, safety considerations lead all agents to active yield. While this reflects legal norms, not all agents minimize risk effectively. For instance, for the standard mode, starting from inf TTC, and risk of 0, and then allows the other vehicle to turn. Yet its approach remains aggressive: by $4\mathrm{s}, TT$C falls to just 0.4s and risk spikes to 0.7, even though no collision occurs. Such behavior is risky and unsuitable for AV deployment. EthicAR w/o PER agent displays a similar pattern. The SACLAG agent demonstrates improved restraint: though ego speed remains relatively high at $4\mathrm{s}, TT$C and risk increase to about 0.3, still yielding more courteously. By contrast, EthicAR maintains the lowest risk level throughout the scenario, embodying responsible and socially considerate driving.
@@ 6 Discussion
In this work, we introduce EthicAR, a hierarchical autonomous driving framework designed to minimize the overall risk to all road users and make ethically responsible decisions. EthicAR not only reacts ethically in high-risk situations but also incorporates ethical considerations into routine driving behavior, thereby proactively avoiding dangerous scenarios. The framework consists of two levels: a decision level and an execution level. The decision level is powered by an LSTM-based SACLag algorithm, enhanced with dynamic PER and trained under ethical cost constraints. The execution level ensures that planned actions are translated into smooth and comfortable vehicle control. To evaluate the contributions of each component, we conduct an ablation study examining the effects of PER, the LSTM architecture, and different cost function designs. The agents are trained using a real-world dataset in simulation, providing realistic scenarios that improve generalization and reliability. For performance evaluation, we test the trained agents on 75 previously unseen real-world scenarios. We analyze the resulting risks for both the ego vehicle and surrounding road users. Our results show that the ethical mode of EthicAR outperforms the other variants, significantly reducing overall risk for all road users compared to versions without cost constraints. Interestingly, although EthicAR is explicitly optimized to reduce risk to all road users, it also achieves lower ego risk than the selfish mode, which considers only the safety of the ego vehicle. This suggests that ethical behavior, inherently designed to protect others, also benefits the ego vehicle, due to the interactive nature of traffic dynamics. Additionally, we analyze four real-world scenarios involving ethical dilemmas. These case studies confirm that EthicAR consistently makes responsible decisions, not only toward other vehicles but also toward VRUs such as cyclists. If stronger prioritization of VRUs is required, this can be readily achieved through simple weight adjustments in the cost function. Looking ahead, extending the framework to incorporate trajectory predictions for pedestrians and cyclists, in addition to vehicles, will enable even more realistic behavior modeling and improve readiness for real-world deployment.